OpenStudy (anonymous):

use distributive property to rewrite expression -2 (9-6)

OpenStudy (anonymous):

The distributive property is: a(b + c) = ab + ac

OpenStudy (anonymous):

So yours would be -2(9-6) = -2(9) - -2(6)

OpenStudy (anonymous):

At least, that is what I believe it is!

OpenStudy (appleboy2):

I'm pretty sure that's correct.